OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality -x - 3 < 13 + 3x. -x - 3 < 13 + 3x -4x < 16 x > -4

jigglypuff314 (jigglypuff314):

mmm "choose" as in that there are options? so are those^ the options or the steps? :/

OpenStudy (anonymous):

steps, want me to put the options?

OpenStudy (anonymous):

-4x - 3 > 13 2x - 3 < 18 -4x - 3 < 13 2x - 3 > 18

OpenStudy (anonymous):

im looking for the second step!

jigglypuff314 (jigglypuff314):

ok yes thanks :) from -x - 3 < 13 + 3x you would try to get all the values that have x with them, to one side... so you would subtract 3x from both sides. that would get you the second step :)

OpenStudy (anonymous):

so -x-<13?

OpenStudy (anonymous):

i appreciate this so much, im a failure when it comes to math. lol its easy when someones going through it with me , otherwise im like a lost puppy.

jigglypuff314 (jigglypuff314):

mmmm more like -x - 3 < 13 + 3x -3x -3x and all you need is more practice :) you've got the basics down

OpenStudy (anonymous):

what would -x - 3x be?

jigglypuff314 (jigglypuff314):

you can think of -x as -1x do you know how to do -1x - 3x ? :)

OpenStudy (anonymous):

-4 - 3 <13?

OpenStudy (anonymous):

so its C ! (:

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

you think i could have help with another?/.\

jigglypuff314 (jigglypuff314):

sure! :)

OpenStudy (anonymous):

Solve for x. -ax + 2b > 8

OpenStudy (anonymous):

your such a sweetiepie (:

jigglypuff314 (jigglypuff314):

-ax + 2b > 8 the "2b" is crowding the x that you want to find, how do you think you can move the "2b" ? :)

OpenStudy (anonymous):

dividing? (taking a wild guess)

jigglypuff314 (jigglypuff314):

lol well "2b" is separated from the x by a " + " and the opposite of " + " would be " - " right? so you would subtract 2b from both sides -ax + 2b > 8 -2b -2b

OpenStudy (anonymous):

so then it's -ax>6?

jigglypuff314 (jigglypuff314):

you wouldn't subtract a number with a variable from a number without a variable so it would just be -ax > 8 - 2b

jigglypuff314 (jigglypuff314):

then, the "-a" is crowding the x it is attached to the x by multiplication so to detach them, you would Divide both sides by Negative a

OpenStudy (anonymous):

so it would be 2b-8/-a

jigglypuff314 (jigglypuff314):

remember to flip the sign when multiplying or dividing by a negative and don't forget the paraenthesis :)

OpenStudy (anonymous):

x < the quantity negative 2 times b plus 8 all over negative a x > the quantity negative 2 times b plus 8 all over negative a x < the quantity 2 times b minus 8 all over negative a x > the quantity negative 2 times b minus 8 all over a

OpenStudy (anonymous):

those are my choices

jigglypuff314 (jigglypuff314):

okay :) after you flip the sign it should become < and from there your guess was correct :)

OpenStudy (anonymous):

thank you! okay last one? i have some of it down i think. i just need guidence through the rest

jigglypuff314 (jigglypuff314):

sure :) tho I might have to go soon :/

OpenStudy (anonymous):

4(2x - 1) - 3 = 9(x - 4) i have 8x-4-3 = 9x -36

OpenStudy (anonymous):

and thats fine, youve helped me enough!

jigglypuff314 (jigglypuff314):

that is correct so far :)

OpenStudy (anonymous):

do i subtract 36?

jigglypuff314 (jigglypuff314):

careful, it is a Negative 36 there so to move it you would add 36 :)

OpenStudy (anonymous):

would i add 36 to 3 on the other side?

jigglypuff314 (jigglypuff314):

on the other side you have -4-3 = -7 so you would then add 36 to -7 so 8x - 7 + 36 = 9x

OpenStudy (anonymous):

8x-1=9x?

jigglypuff314 (jigglypuff314):

??? 36 - 7 = ?

OpenStudy (anonymous):

oh so 8x-29?

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

then subtract 8x from both sides

OpenStudy (anonymous):

x=21?

OpenStudy (anonymous):

wait im confused... thats not one of my choices

jigglypuff314 (jigglypuff314):

be careful to not subtract a number that has x attached to it with a number that doesn't have x attached like 2x - 3 would just be 2x - 3 so 8x + 29 = 9x -8x -8x

OpenStudy (anonymous):

29=1x

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

so my answer is 29?

jigglypuff314 (jigglypuff314):

I think so :)

OpenStudy (anonymous):

Thank you again for all your help! i hope you have a fantastic day! (:

jigglypuff314 (jigglypuff314):

glad I could help :) best of luck on your studies!

OpenStudy (anonymous):

ughh thank you! lol i dont know how im going to finish this D:

jigglypuff314 (jigglypuff314):

you can tag people to your question by putting a @ in front of their name I might come back later, but if you still need more help some good people are whpalmer4 tHe_FiZiCx99 sourwing Loser66 johnweldon1993 jdoe0001 eliassaab just try not to tag them all a once or anything, since they've probably got a whole line of people trying to get them smart people to answer their question. Good Luck! :)